OpenStudy (anonymous):

Which number is a rational number? A. √21 B. √36 C. √50 D. √63 Help please? Will give a medal. :-)

OpenStudy (welshfella):

look for one that has a whole number as a square root

OpenStudy (anonymous):

a rational number means it can be writen as n/m where n and m are integers, m not equal to 0

OpenStudy (ybarrap):

$$ \sqrt{21}=\sqrt{3\times 7}=\sqrt 3\sqrt 7\\ \sqrt{36}=\sqrt{6^2}=6\\ \sqrt{50}=\sqrt{5\times 10}=\sqrt 5\sqrt {10}=\sqrt 5\sqrt {2\times 5}=\sqrt 5\sqrt {2}\sqrt{5}=5\times \sqrt 2\\ \sqrt{63}=\sqrt{7\times 9}=3\sqrt 7 $$ All the ones with square roots above, which can not be expressed as an integer or a fraction with integers (i.e numerator and denominators as integers) are NOT rational.